OpenStudy (anonymous):

Solve for w -3w – 25 = -5(w + 7) Simplify your answer as much as possible

OpenStudy (anonymous):

This is easy. first distribute. -3w-25=-5w-35. what is the next step??

OpenStudy (anonymous):

-3w-25=-5w-7 +25 +25 -3w =-5w+18 +5w +5w 2w/2=18/2 w=9

OpenStudy (anonymous):

5 = w, if I'm not mistaken. :)

OpenStudy (anonymous):

it should be 9

OpenStudy (anonymous):

its -5

OpenStudy (anonymous):

Wait @Niall is right i messed up

OpenStudy (anonymous):

Its -5, I didn't push the - sign hard enough the first time. :)